A 3.82 g sample of magnesium nitride is reacted with 7.73 g of water.
Mg3N2 + 3H2O → 2NH3 + 3MgO
The yield of MgO is 3.60 g. The percentage yield of the reaction is ________%
1
Expert's answer
2017-02-27T15:30:05-0500
Mg 3 N 2 + 3H 2 O → 2NH 3 + 3MgO

n(Mg 3 N 2 ) = 3.82/100 = 0.0382 mol

n(H 2 O) = 7.73/18 = 0.4294 mol

Mg 3 N 2 ‒ limiting reagent

100 g Mg 3 N 2 ‒ (3 x 40) g MgO

3.82 g Mg 3 N 2 ‒ x g MgO

x = 4.584 g MgO

η = 3.6/4.584 x 100 = 78.5 %

Answer: η = 78.5 %